A solution of cuso4 is electrolysed for 10min by passing a current of 1amp.caculate the weight of cu deposited at electrode

Grade:12

1 Answers

PRAJWAL
6 years ago
10 min =600 sec, I=1 A Cu2+ there fore for the deposition of molar mass of copper 2*96500C of charge is required Q=It=1*600=600c
accordingly-2*96500    60g
                       600          ?
600*60/2*96500 is the mass of copper deposited by passing 1 amp of current for 10 min